OpenStudy (anonymous):

Your probability of getting the flu this year is .15. The probability that you will get the flu and run a high fever is .05. Suppose you come down with the flu. What is the probability that you will get the high fever, given that you have the flu?

OpenStudy (anonymous):

let the probablility of getting the flu be p, probability of getting a fever be q. We know p = .15, and we know p*q = .05

OpenStudy (anonymous):

so q = .05/p = (.05)/(.15) = 1/3

OpenStudy (anonymous):

or .33333.... or 33%

OpenStudy (anonymous):

\[\sqrt{\log x} = \log \sqrt{x}, can any one solve for x\]

OpenStudy (anonymous):

thank you :D